Medal Corporation uses direct labor hours to allocate factory overhead. During the period, the company produced 5,100 units and incurred variable factory overbead costs of $21,000, Use the information below to determine the following (round dollars to the nearest cent Medal Corporation Factory Overhead Cost Budget For the three months ending March 31, 2015 100% Percent of normalc Units produced Direct labor hours (0.75 hr. per unit) 110% 4.500 500 125 3.750 udgeted Variable cost: Indirect Indirect materials 8,100 9,000 ,000 3.600 18,450 12,000 20,000 tilities Total variable cost Supervisor salaries Total fixed costs 500 Fixed costs 12 20.0 Total overhead cost a. Variable factory overhead rate b. Variable factory overhead controllable variance c. Is the variable factory overhead controllable variance favorable or unfavorable? d. Fixed factory overhead rate e. Fixed factory overhead volume variance . Is the fixed factory overhead volume variance favorable or unfavorable? g. Determine the total factory overhead cost variance by comparing actual and applied factory overhead. Is the variance favorable or unfavorable?
Answer #1

Solution a:

Variable factory overhead rate = Budgeted variable overhead / Budgeted direct labor hours = $18,450 / 3375 = $5.47 per hour

Solution b:

Variable factory overhead controllable variance = Standard variable overhead cost - Actual variable overhead cost

= (5100*0.75*$5.466666) - $21,000 = $90 U

Solution c:

Variable factory overhead controllable variance in unfavorable.

Solution d:

Fixed factory overhead rate = Budgeted fixed overhead at normal capacity / Budgeted labor hours at normal capacity

= $20,000 / 3750 = $5.33 per labor hour

Solution e:

Fixed overhead applied = Standard hours * Fixed factory overhead rate = 5100*0.75*$5.3333333 = $20,400

Fixed factory overhead volume variance = Fixed overhead applied - Budgeted fixed overhead

= $20,400 - $20,000 = $400 F

Solution f:

Fixed factory overhead volume variance is favorable.

Solution g:

Total overhead applied = Standard variable overhead + Fixed overhead applied

= (5100*0.75*$5.466666) + $20,400 = $41,310

Actual overhead incurred = $21,000 + $20,000 = $41,000

Total factory overhead cost variance = Total overhead applied - Actual overhead applied = $41,310 - $41,000 = $310 F

The variance is favorable.